OpenStudy (anonymous):

Find the point (0, b) on the y-axis that is equidistant from the points (4, 4) and (4, -3).

OpenStudy (freckles):

so you have an equation you want the distance from (0,b) to (4,4) to be the same as the distance from (0,b) to (4,-3)

OpenStudy (freckles):

use distance formula twice once for both sets of points then equate and solve for b

OpenStudy (anonymous):

well I got -1/14, but its saying its wrong

OpenStudy (freckles):

so can I ask what equation you solved? did you solve the one I asked you to?

OpenStudy (anonymous):

yes I used two distance formula for (0,b)(4,4) and (0,b)(4,-3) and i got 24-8b+b^2=25+6b+b^2 which i got 1/14

OpenStudy (freckles):

\[\sqrt{(0-4)^2+(b-4)^2}=\sqrt{(0-4)^2+(b-(-3))^2} \\ \\ \text{ square both sides } \\ (0-4)^2+(b-4)^2=(0-4)^2+(b-(-3))^2\] now 0-4 is -4 and (-4)^2 =16 also b-(-3) is the same as b+3 so you have: \[16+(b-4)^2=16+(b+3)^2 \\ \text{ subtract 16 on both sides } (b-4)^2=(b+3)^2\] is this what you try to solve ?

OpenStudy (freckles):

this means you either have b-4=b+3 or b-4=-(b+3) see which one of these equations gives you a solution

OpenStudy (freckles):

I'm gonna check on my croissant dough unless you have a question

OpenStudy (anonymous):

im just trying it out to see, i think ill be fine

OpenStudy (freckles):

ok well I will be back in a few if not

OpenStudy (anonymous):

ah i got it, thank you so much

OpenStudy (freckles):

awesome